OpenStudy (ivancsc1996):

Why do we write that the inverse function of sin is:

OpenStudy (ivancsc1996):

\[\sin ^{-1}\]since the inverse function of sin isn't \[\frac{ 1 }{ \sin }\]

OpenStudy (anonymous):

the exponent of \(-1\) means the inverse of the function under composition, not the inverse of the number under multiplication

OpenStudy (anonymous):

with numbers, \(a^{-1}=\frac{1}{a}\) because \[a\times a^{-1}=1\] and \(1\) is the identity under multiplication

OpenStudy (anonymous):

the inverse of a function \(f^{-1}\) is the function that, when composed with \(f\) gives the identity function \(I(x)=x\)

OpenStudy (anonymous):

that is, \[f^{-1}\circ f=I\]

OpenStudy (anonymous):

means give the angle that has x for its sin that's not really true when you go further into trigonometry.

OpenStudy (anonymous):

"means give the angle that has x for its sin"*

OpenStudy (anonymous):

the operation is composition, not multiplication, and the identity is the identity function, not the number 1

OpenStudy (anonymous):

the notation itself has nothing to do with trigonometry, it is the notation used for all functions for example if \(f(x)=2x-1\) then \(f^{-1}(x)=\frac{x+1}{2}\) not \(\frac{1}{2x-1}\)

OpenStudy (anonymous):

if you really have trouble using sin^-1, use arcsin you will be well-understood